# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

### v2

@@ II.2 Disentangling Fitzpatrick’s expressions
The poloidal magnetic flux is the flux enclosed in the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$.
@@
A fundamental mathematical point that Fitzpatrick missed is that the toroidal magnetic flux is defined by an inside-outwards area integral from the magnetic axis while the poloidal flux is defined by an outward-inwards integral towards the magnetic axis.
Fitzpatrick’s condition for a sufficiently slow ramp up or down to avoid a disruption, his Equation (42), is equivalent to
@@
His condition can be expressed as a condition on the loop voltage at the plasma edge, $V_{\ell}^{edge}=(d\psi_{p}/dt)_{edge}$, which is at $r=a$. The poloidal flux at the plasma edge is given by
@@
During a ramp up, $d\psi_{p}^{in}/dt>0$, Fitzpatrick’s condition, Equation (20), for avoiding a disruption is $V_{\ell}^{edge}>0$, which is a necessary condition for any ramp up but not a sufficient condition for avoiding a disruption. During a ramp-down, $d\psi_{p}^{in}/dt<0$, Fitzpatrick’s condition for avoiding a disruption is $V_{\ell}^{edge}<2V_{\ell}^{ax}$, which also not a sufficient condition for avoiding a disruption. What is required is that the current density profile not be disruptive as discussed in Subsection II.4.
II.3 Fitzpatrick check for non-disruptive profiles
Although Fitzpatrick gave an expression for the profile of perpendicular energy diffusivity of some generality in his Equation (88), he actually determined the current profile that would be obtained for only one case, a radially constant energy-diffusivity profile. He then checked that the resulting current profile is not disruptive using the internal inductance $\ell_{i}$, Fitzpatrick’s Equation (90), and the edge safety factor $q_{a}$ of the 1987 Cheng, Furth, and Boozer study MHD stab in a cylindrical approximation of disruptions in TFTR.
What Fitzpatrick did not do is determine the range of profiles of his energy diffusivity that give non-disruptive current profiles even by the Cheng et al criterion that he was using. Showing stability for one profile of the heat diffusivity is certainly no guide for determining the danger of disruptions.
In his response to an earlier version of this Comment, Fitzpatrick stated Fitzpatrick:2026A : “The transport of electron energy in tokamak plasma has been the subject of extensive research for the last 50 years. There is overwhelming evidence that such transport is diffusive in nature.”
Despite the dogmatism of Fitzpatrick’s statement, the general view among tokamak experts is that energy diffusivity is not always the correct way to characterize energy transport. The temperature profiles can be given rigidity by microturbulence, Holland:2021 ; Turb-profiles:2025 . The study by N.T. Howard et al of microturbulent transport in ITER Turb-profiles:2025 has sufficient information to calculate the loop voltage and determine its spatial constancy. This would determine the consistency of their assumed current profile $I(\psi_{t})$. Much work remains to establish the consistency of current profiles with both thermal transport and stability against disruptions in full ramp-up, flattop, and ramp-down scenarios even as the nuclear power production turns on and off.
@@ II.4 Actual condition for disruption avoidance
Not only did Fitzpatrick ignore the simple statements in my article about the poloidal flux and its evolution, he also ignored the statement in the abstract in my paper about the important result: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
Fitzpatrick could not examine the correctness of this statement for two reasons. First, by ignoring $\psi_{p}^{ex}$ he ignored most of the poloidal flux produced by the plasma current. Second, by not considering the range of profiles of current density that are consistent with stability by the Cheng-Furth-Boozer criterion MHD stab , which Fitzpatrick used in his paper, he could not estimate what range of poloidal fluxes are stable to disruptions.
Only a small fractional change in the poloidal flux is required to move a tokamak equilibrium that is stable to disruptions into a disruptive state. Consequently, a very careful control of the poloidal flux evolution is required during a plasma shut down—all the poloidal flux produced by the plasma current $\Psi_{p}=-(\psi_{p}^{in}+\psi_{p}^{ex})$ must be removed. The length of the timescale for $\Psi_{p}$ removal is so long that the current profile within the plasma has time to relax to nearly its steady-state limit, which is a spatially constant loop voltage $V_{\ell}$ across the plasma. Ignoring current drive and the bootstrap current, as Fitzpatrick does, the net toroidal current density on each magnetic surface, which is denoted by its enclosed toroidal flux $\psi_{t}$, is
@@
where $2\pi R$ is the average distance required for a field line to make a toroidal circuit, $Z_{eff}$ is the correction to the resistivity produced by impurities, and $T_{e}$ is the electron temperature. Neoclassical effects to the resistivity $\eta$ can be easily included. The difficulty of obtaining disruption-free shutdowns in JET was discussed in Sozzi:2020 . This was thought largely due to impurity issues.
@@
Although not emphasized in his paper, Fitzpatrick assumed a spatially constant loop voltage in the plasma region in both ramp-up and ramp-down of the plasma current. The primary point in my Physics of Plasmas submission was the spatial constancy of the loop voltage. A major difference in calculations is that my article considered a number of current profiles that ranged over the stable region of the Cheng et al MHD stab paper and found the poloidal flux in the plasma changed by relatively small amount over that range of current profiles. Fitzpatrick considered only one current profile, which happened to be stable, but this is not a credible argument for the robust stability against disruptions for all realistic current profiles.

### v3

@@ II.2 Disentangling the expressions used in Fitzpatrick:2026
The poloidal magnetic flux is the flux that goes down through the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$.
@@
A fundamental mathematical point that was missed in Fitzpatrick:2026 is that the toroidal magnetic flux is defined by an inside-outwards area integral from the magnetic axis while the poloidal flux is defined by an outward-inwards integral towards the magnetic axis from the $R=0$ axis of the torus.
The necessary condition in Fitzpatrick:2026 for a sufficiently slow ramp up or down to avoid a disruption is given by Equation (42), which is equivalent to
@@
This condition can be expressed as a condition on the loop voltage at the plasma edge, $V_{\ell}^{edge}=(d\psi_{p}/dt)_{edge}$, which is at $r=a$. The poloidal flux at the plasma edge is given by
@@
During a ramp up, $d\psi_{p}^{in}/dt>0$, the necessary condition Equation (20), which is a rewrite of Equation (42) in Fitzpatrick:2026 , for avoiding a disruption is $V_{\ell}^{edge}>0$, which is a necessary condition for any ramp up but not a sufficient condition for avoiding a disruption. During a ramp-down, $d\psi_{p}^{in}/dt<0$, the necessary condition of Fitzpatrick:2026 for avoiding a disruption is $V_{\ell}^{edge}<2V_{\ell}^{ax}$, which is also not a sufficient condition for avoiding a disruption. What is required is that the current density profile not be disruptive as discussed in Subsection II.4.
II.3 Non-disruptivity of the current profile of Fitzpatrick:2026
Although the paper Fitzpatrick:2026 gave an expression for the profile of perpendicular energy diffusivity of some generality, Equation (88), the paper actually determined the current profile that would be obtained for only one case, a radially constant energy-diffusivity profile. The resulting current profile is not disruptive using the internal inductance $\ell_{i}$, Equation (90) in Fitzpatrick:2026 , and the edge safety factor $q_{a}$ based on the 1987 Cheng, Furth, and Boozer study MHD stab in a cylindrical approximation of disruptions in TFTR. In the paper Fitzpatrick:2026 , the only source of heating was Ohmic and the only source of cooling was the thermal transport coefficient, which obeyed a strong assumption—a fixed profile in time that was spatially constant. In reality, fusion power is centrally concentrated and quickly ceases during shutdown. The radiated power from impurities has a complicated dependence on temperature, and impurities modify the resistivity profile through the $Z_{eff}$ dependence of the resistivity $\eta$.
What was not done in Fitzpatrick:2026 is to determine the range of profiles of the energy diffusivity that give non-disruptive current profiles by the Cheng et al criterion—even in the approximation of ignoring the subtleties of heating and radiative cooling. Showing stability for one time-independent thermal-diffusivity profile, which was a spatial constant, is not a reliable guide for determining the danger of disruptions.
In a response to an earlier version of this Comment, Fitzpatrick Fitzpatrick:2026A stated: “The transport of electron energy in tokamak plasma has been the subject of extensive research for the last 50 years. There is overwhelming evidence that such transport is diffusive in nature.” Despite the certainty of this statement, the general view among tokamak experts is that energy diffusivity is not always the correct way to characterize energy transport—even when radiative losses are ignored. The temperature profiles can be given rigidity by microturbulence, Holland:2021 ; Turb-profiles:2025 . The study by N.T. Howard et al of microturbulent transport in ITER Turb-profiles:2025 has sufficient information to calculate the loop voltage and determine its spatial constancy. This would determine the consistency of their assumed current profile $I(\psi_{t})$. Much work remains to establish the consistency of current profiles with both thermal transport and stability against disruptions in full ramp-up, flattop, and ramp-down scenarios even as the nuclear power production turns on and off and as impurities gives complicated power losses.
@@ II.4 Actual condition for disruption avoidance
Not only did the paper Fitzpatrick:2026 ignore the simple statements in the arXiv article Boozer:flux about the poloidal flux and its evolution, the paper also ignored the statement in the abstract Boozer:flux about an important result: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
The paper Fitzpatrick:2026 could not examine the correctness of this statement for two reasons. First by ignoring $\psi_{p}^{ex}$, it ignored most of the poloidal flux produced by the plasma current. Second by not considering the range of profiles of current density that are consistent with stability by the Cheng-Furth-Boozer criterion MHD stab , the paper Fitzpatrick:2026 could not estimate what range of poloidal fluxes produced by the plasma current are stable to disruptions.
Only a small fractional change in the poloidal flux is required to move a tokamak equilibrium that is stable to disruptions into a disruptive state. Consequently, a very careful control of the poloidal flux evolution is required during a plasma shut down—all the poloidal flux produced by the plasma current $\Psi_{p}=-(\psi_{p}^{in}+\psi_{p}^{ex})$ must be removed. The length of the timescale for $\Psi_{p}$ removal is so long that the current profile within the plasma has time to relax to nearly its steady-state limit, which is a spatially constant loop voltage $V_{\ell}$ across the plasma. Ignoring current drive and the bootstrap current, as the paper Fitzpatrick:2026 does, the net toroidal current density on each magnetic surface, which is denoted by its enclosed toroidal flux $\psi_{t}$, is
@@
where $2\pi R$ is the average distance required for a field line to make a toroidal circuit, $Z_{eff}$ is the correction to the resistivity produced by impurities, and $T_{e}$ is the electron temperature. Neoclassical effects to the resistivity $\eta$ can be easily included. Both the $T_{e}$ and $Z_{eff}$ profiles generally evolve as the plasma evolves and are not independent of time as assumed in Fitzpatrick:2026 . The difficulty of obtaining disruption-free shutdowns in JET was discussed in Sozzi:2020 . This was thought largely due to impurity issues.
@@
The paper Fitzpatrick:2026 assumed a poloidal field profile that was essentially independent of time, which implies a spatially constant loop voltage in the plasma region. However, a spatially constant loop voltage does not imply a time independent poloidal field profile when the electron temperature and the $Z_{eff}$ profiles are time dependent.
The primary point in the Physics of Plasmas submission, which was the September 9, 2025 version of Boozer:flux , was the spatial constancy of the loop voltage. A major difference in calculations is that the arXiv article Boozer:flux considered a number of current profiles that ranged over the stable region of the Cheng et al MHD stab paper and found the poloidal flux in the plasma changed by relatively small amount over that range of current profiles.
The Nuclear Fusion paper Fitzpatrick:2026 considered only one current profile, which happened to be stable, but this is not a credible argument for the robust stability against disruptions for all realistic current profiles.